### v1

@@ 3 Results
We developed subject-specific gyrencephalic and lissencephalic FE head models of 18 individuals from their MRI scans. Each model consisted between $\sim 1.7-2.7\times 10^{6}$ reduced integration hexahedral elements (C3D8R). We simulated a concussive blunt impact on each model through rigid body rotation of the skull (Fig 1d). The simulations were performed on Abaqus/Explicit and each simulation took 10-14 hours using 64 cores. In this section, we describe the injury metrics (defined in Section 2.3) obtained from these simulations to understand the effect of modeling cortical folds.
3.1 Injury risk in Gyrencephalic vs Lissencephalic Models: A Representative Case
This section analyzes the evolution of strain and strain rate spatial distributions during axial head rotation in a representative subject (18F) to identify the key differences in trends between the gyrencephalic and lissencephalic models (Figure 2). The spatial distributions are shown across an axial slice passing through the anterior and posterior horns of the lateral ventricles, parallel to the horizontal plane for the two models.
@@ 3.1.1 Brain tissue strain
We found some similarities in the overall brain deformation wave motions in the two models. The maximum principal strain (MPS) distributions at different times show that the strain wave originates at the outer surface of the brain (cortex) and travels to the center of the brain in both the gyrencephalic and lissencephalic models (Figure 2a). The strain dissipates as the wave travels inwards, resulting in higher strain concentration in the cerebral cortex in both models, before localizing near ventricles (near the anterior horns of the lateral ventricles) (Figure 2a).
@@
The peak maximum principal strain (PMPS) at a point over the course of the simulation enables high-strain region comparison between two models by removing the consideration of the time of peak strain. The PMPS distribution shows that the lissencephalic model experienced high strains near the outer surface of the brain and the gyrencephalic model near the ventricles and deeper regions of the brain in addition to the outer cortical surface (Figure 2c). The CSDM15, which is the region above an injury threshold, shows a similar trend in the damaged regions (Figure 2d).
@@ 3.1.2 Brain tissue strain rate
Figure 2: a) Maximum principal strain (MPS) contour plots of a representative subject at different times near MPS95 peak from the lissencephalic and the gyrencephalic models, b) the peak MPS (PMPS) experienced by the element over the entire duration of the head acceleration event, c) damaged element shown in red after the MPS at its location exceeded 0.15. These plots highlight the spatial variations arising due to the cortical folds.
3.2 Inter-Subject Variability in Strain Distribution Differences Between Gyrencephalic and Lissencephalic Models
@@
3.2.1 Axial Head Rotation
Under the axial head rotations, both the gyrencephalic and lissencephalic models of all subjects experienced strain concentrations (PMPS) on the cortical surface and around the ventricles, notably between the anterior horns of the ventricles and the lateral (Sylvian) sulcus (Figure 3a). The strain concentrations near the anterior horns of the lateral ventricles in lissencephalic models were lower magnitudes compared to the gyrencephalic models.
The outer cortical regions under higher strain concentrations are non-uniform and variable among the subjects in the gyrencephalic models. On the other hand, the high-strain outer cortical regions in lissencephalic models are consistently present along the lateral regions (the smaller axis of the cross-section ellipse), where the radius of curvature is high. The strain concentration near the ventricles becomes more prominent with increasing moment of inertia (MoI) about the axial axis, in both models (left to right Figure 3a).
@@ 3.2.1 Axial Head Rotation
The DICE similarity coefficient (mean $\pm$ SD) between gyrencephalic and lissencephalic models ranged between $0.17\pm 0.07$ for PMPS distributions, and between $0.60\pm 0.10$ for CSDM under the axial loading. The CSDM15 DICE is much higher than PMPS, since CSDM15 DICE essentially discretizes the PMPS distribution in steps of 1.0, while PMPS DICE considers similarity in steps of 0.025.
3.2.2 Sagittal Head Rotation
Similarly, Figure 3b shows the PMPS contours across a sagittal plane through the temporal lobe for different subjects under the sagittal head rotations. The gyrencephalic models show high strain regions on the outer cortical surface and near the sulci in the frontal and temporal lobes, whereas the lissencephalic models only show high strains on the cortical surface. The DICE similarity coefficients for the whole brain regions ranged between $0.19\pm 0.08$ for PMPS, and $0.52\pm 0.10$ for CSDM for the sagittal rotations.
3.2.3 Coronal Head Rotation
Similar trends were also seen under coronal head rotations in Figure 3c, where the strain contours for representative subjects are shown across a coronal slice through the frontal and temporal lobes. The gyrencephalic models show high strain regions on the outer cortical surface and around the Sylvian sulcus in the frontal and temporal lobes. The lissencephalic models only capture the strain concentration on the outer cortical surface. The DICE similarity coefficients for the whole brain regions ranged between $0.22\pm 0.10$ for PMPS, and $0.59\pm 0.08$ for CSDM for the sagittal rotations.
These results show that modeling cortical folds consistently predicts higher strain concentrations between the ventricles and the deeper sulci, such as the lateral (Sylvian sulci), across all subjects and rotation directions, which couldn’t be captured in lissencephalic models.
@@ 3.2.3 Coronal Head Rotation
Figure 3: Peak maximum principal strain (MPS) contour plots of multiple representative subjects in the lissencephalic and the gyrencephalic models under rotational acceleration about each principal axis: a) axial, b) sagittal, and c) coronal. These PMPS distributions capture the interaction of cortical folds with ventricles to result in different strain localization, especially around deeper sulci, such as the lateral (Sylvian) sulcus.
@@ 3.3 Effect of cortical folds on peak injury metrics
The time history plots of injury metrics, including MPS95, MPSR95, and CSDM, for all the subjects and under the three loading directions show that the lissencephalic models consistently under-predicted all injury metrics (Figure 4). The time to peak MPS95 and MPSR95 is also lower in the lissencephalic models for all loading directions, and the CSDM15 is initially higher in the lissencephalic models, before being surpassed by the gyrencephalic models (Figure 4).
@@
Figure 5: Bland-Altman analysis showing the mean bias and the limits of agreement (mean±1.96×SD) between the gyrencephalic and lissencephalic peak injury metrics: a) MPS95, b) MPSR95, and c) CSDM15. These trends show the gyrencephalic models overpredict the injury metrics compared to the lissencephalic models with high statistical significance.
@@ 3.4 Effect of modeling cortical folds on different regions of interest
3.4.1 Cerebral Gray Matter
The cerebral gray matter is geometrically different between the gyrencephalic and the lissencephalic models. The region experienced overall higher MPS95, MPSR95, and CSDM15 in the gyrencephalic models than the lissencephalic ones for all three loading directions ($\textit{p}<0.005$). The Bland-Altman analysis provided the bias (mean $\pm$ std. dev.) in peak MPS95 to be $0.086\pm 0.057$ (statistically significant, $p<0.005$), in peak MPSR95 to be $15.4\pm 10.7$ s-1 ($p<0.005$), and in CSDM15 to be $11\pm 8.2$ ($p<0.005$). (Figure 6a). The bias values translate to $17.7\pm 11.8\%$, $16.0\pm 11.2\%$, and $12.0\pm 9.0\%$ of the maximum MPS95, MPSR95, and CSDM15 values in the gyrencephalic models gray matter, respectively. No proportional bias was observed in any injury metric.
3.4.2 Cerebral White Matter
The cerebral white matter is also geometrically different between the gyrencephalic and the lissencephalic models. The region experienced overall higher MPS95 and MPSR95 in the gyrencephalic models than the lissencephalic ones under the sagittal and coronal rotation directions, but consistently lower MPS95 and MPSR95 under the axial direction. The Bland-Altman analysis provided the bias (mean $\pm$ std. dev.) in peak MPS95 to be $0.006\pm 0.035$ (statistically insignificant, $p>0.05$), in peak MPSR95 to be $2.5\pm 5.3$ s-1 ($p>0.05$), and in CSDM15 to be $3\pm 11.73$ ($p>0.05$). (Figure 6a). The normalized bias values are $1.8\pm 9.9\%$, $3.4\pm 7.2\%$, and $7.2\pm 3.0\%$ in MPS95, MPSR95, and CSDM15, respectively, highlighting the small difference and high variability across individuals and loading directions. No proportional bias was observed in any injury metric.
We see that the axial loading condition causes higher strains in the lissencephalic models, which could be from differences in modeling the gray and white matter in the two models. Based on the size of high strain concentration on the lissencephalic cortical surface, lissencephalic white matter might experience higher strain, whereas the same depth from the cortical surface falls in the gray matter in the gyrencephalic model.
@@ 3.4.3 Cerebellum
The cerebellum is geometrically identical between the gyrencephalic and the lissencephalic models and away from the cortical folds separated by tentorium. The region experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions ($\textit{p}<0.005$).
@@
The bias (mean $\pm$ std. dev.) in peak MPS95 was found to be $0.032\pm 0.041$ (statistically significant, $p<0.005$), in peak MPSR95 to be $6.36\pm 7.88$ s-1 ($p<0.005$), and in CSDM15 to be $11\pm 11.22$ ($p<0.005$) (Figure 6c). The normalized bias values are $7.6\pm 10.0\%$, $8.5\pm 8.5\%$ and $11.59\pm 12.4\%$, in MPS95, MPSR95, and CSDM15, respectively. We observe proportional bias in all injury metrics, with the bias increasing with the magnitude of the injury metric.
3.4.4 Deep Brain Regions
@@
The deep brain regions (Section 2.1) is geometrically identical between the gyrencephalic and the lissencephalic models and separated from the cortical folds through white matter around it. The region also experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions ($\textit{p}<0.005$).
@@
We found the mean bias ($\pm$ std. dev.) in peak MPS95 was found to be $0.056\pm 0.038$ (statistically significant, $p<0.005$), in peak MPSR95 to be $12.89\pm 8.18$ s-1($p<0.005$), and in CSDM15 to be $16\pm 11.22$ ($p<0.005$) (Figure 6d). The normalized bias values are $13.1\pm 8.5\%$, $14.9\pm 8.4\%$ and $19.9\pm 14.2\%$, in MPS95, MPSR95, and CSDM15, respectively. A proportional bias was observed again in all injury metrics, with the bias increasing with the magnitude of the injury metric.
3.4.5 Corpus Callosum
The corpus callosum is geometrically the same between the gyrencephalic and the lissencephalic models, although is located close to the gyri on the medial axis. It experienced one of the highest peak values for all injury metrics and also the highest biases between the two models ($40\%$) across all loading directions. It experienced higher strain in the gyrencephalic than in the lissencephalic models for all three loading directions ($\textit{p}<0.005$). (Figure 6e)

### v2

@@ 3 Results
We developed subject-specific gyrencephalic and lissencephalic FE head models of 18 individuals from their MRI scans. Each model consisted between $\sim 1.7-2.7\times 10^{6}$ voxel hexahedral elements. We simulated a concussive blunt impact on each model through rigid body rotation of the skull (Fig 1d). The simulations were performed on Abaqus/Explicit and each simulation took 10-14 hours using 64 cores. In this section, we describe the tissue-level injury metrics (Section 2.3) obtained from these simulations to understand the effect of modeling cortical folds.
3.1 Effect of modeling cortical folds on injury metrics: A representative case
This section analyzes the spatial evolution of the strain and strain rates during axial head rotation in a representative subject (18F) to identify the key differences in trends between the gyrencephalic and lissencephalic models (Figure 2). The spatial distributions are shown across an axial slice passing through the anterior and posterior horns of the lateral ventricles, parallel to the horizontal plane for the two models.
@@ 3.1.1 Brain tissue strain
We found some similarities in the overall brain strain waves in the two models. The MPS distributions at different times show that the strain begins to increase at the outer surface of the brain (cortical surface) and travels to the center of the brain in both the gyrencephalic and lissencephalic models (Figure 2a). We see higher strain concentration in the cerebral cortex in both models, although to different degrees (Figure 2a). Since a head acceleration event is modeled through providing the skull with a rigid body acceleration, a strain wave originates at the skull and travels inwards to the brain through the meninges and CSF. While the spherical convergence of the strain wave tends to increase the strain magnitude as the wave travels inwards, the high viscosity of the brain tissue causes dissipation of the strain magnitude as the wave travels inwards, and therefore results in higher strains on the cortical surface (Figure 2a-f) [44, 45].
We observed differences between the two models in terms of the location of high-strain regions on the cortical surface. The lissencephalic model experiences higher strain on the lateral regions, which decreases gradually moving towards the falx cerebri along the cortical surface. Whereas, the high strain regions in the gyrencephalic model are non-uniform. In the gyrencephalic model, the space between sulci is filled with CSF, which is significantly softer than the brain tissue material occupying the space in the lissencephalic model [29]. Therefore, modeling cortical folds directly impacts the effective stiffness of the skull-brain interface, which plays a significant role in the strains experienced by the brain tissue. The thickness of the CSF layer varies across gyri, contributing to the heterogeneity of the high-strain region on the cortical surface. Also, the size and orientation of different sulci vary, resulting in different moments of inertia about the rotation axis, which can also affect the strain distributions. As the strain wave travels inwards, the geometric discontinuity from sulci (high curvature voids) causes higher strain at the tip of the sulci as compared to the adjacent tissue lying at the same depth from the surface of the gyri.
@@
Overall, the lissencephalic model experienced high strains near the outer surface of the brain, whereas the gyrencephalic model experienced its highest strains near the ventricles and deeper regions of the brain in addition to the outer cortical surface (Figure 2c), causing the time to peak MPS95 to be lower in the lissencephalic model ($t_{L}=10$ms) than in the gyrencephalic model ($t_{G}=12$ms). The CSDM15, which is the region above an injury threshold, shows a similar trend in the damaged regions (Figure 2d). The spatial CSDM15 distribution has a higher DICE similarity coefficient of 0.67 between the models.
@@ 3.1.2 Brain tissue strain rate
The MPSR distributions show similar trends as MPS while comparing the gyrencephalic and the lissencephalic models (Figure 2d-f). The MPSR distributions show the strain rate wave propagating from the outer cortical surface to the center of the brain, while dissipating slowly (Figure 2d). The lissencephalic model shows higher MPSR and CSRDM50 on the cortical surface than the gyrencephalic model distributions (Figure 2e-f). On the other hand, the highest MPSR and CSRDM50 in the gyrencephalic model are seen near the center of the brain around the ventricles, due to the interaction between the regions of strain concentrations at the boundary of voids created by the sulci and the ventricles, as explained in Section 3.1.1.
@@
Figure 2: a) Maximum principal strain (MPS) contour plots of a representative subject at different times near MPS95 peak from the lissencephalic and the gyrencephalic models, b) the peak MPS (MPS) experienced by the element over the entire duration of the head acceleration event, c) damaged element shown in red after the MPS at its location exceeded 0.15. These plots highlight the spatial variations arising due to the cortical folds.
3.2 Inter-subject variability in injury risk due to modeling cortical folds
@@
3.3 Effect on spatial strain distributions
Both the gyrencephalic and lissencephalic models of all subjects experienced higher strains on the outer cortical surface under all rotation directions. The high-strain regions on the outer cortical surface in lissencephalic models are present consistently along the smaller axis of the cross-section ellipse, where the radius of curvature is high, which allows the brain tissue to deform unconstrained under all loading directions. These high-strain regions on the cortical surface include i) the lateral regions away from the falx cerebri under the axial rotations (Figure 3a); ii) regions along the tentorium cerebeli, along the convex curvature regions between the frontal and the temporal lobe, and along the top cortical surface under the sagittal rotations (Figure 3b); and iii) on the lateral cortical surface away from the falx cerebri and the convex curvature region near the inferior cortical surface under the coronal rotations (Figure 3c). (These regions correspond to high-strain regions in solids with different cross-sections under rotation, Supplementary Section S5). In the previous section, we saw that the skull-brain interface plays a significant role in governing the brain deformation, and the gyrencephalic model captures the heterogeneous skull-brain interface. Therefore, the individual variations in the cortical folds make the high-strain regions in the gyrencephalic model heterogeneous along the cortical surface as well as across different individuals. The CSF thickness between the brain and the skull also has a cushioning effect, dissipating the strain reaching the brain tissue, (Supplementary Section S4), also contributing to the hetergeneity in the brain tissue strain along the cortical surface.
In the inner regions of the brain, we observe more strain concentrations in the gyrencephalic model across all individuals and loading directions. Under axial rotations, the high strain concentration regions between the anterior horns of the lateral ventricles and the lateral (Sylvian) sulcus in gyrencephalic models become increasingly larger with increasing head size (moment of inertia, MoI about the axial axis) as compared to the lissencephalic models. Under sagittal rotations, the gyrencephalic models experience higher strains near the posterior horns of the lateral ventricles and the sucli in the frontal and temporal lobes. Similarly, under the coronal rotations as well, the gyrencephalic models show high strain regions on the outer cortical surface and around the Sylvian sulcus in the frontal and temporal lobes. The individual differences in the cortical fold geometry cause higher individual variations in the gyrencephalic model. In the previous section, we also saw that the distance between the ventricles and sulci affects the strain concentration in the gyrencephalic model, and consequently, the individual variations contribute to higher variations in the high-strain regions in the gyrencephalic models across individuals [46]. These differences across individuals are not pronounced in the lissencephalic models.
@@ 3.3 Effect on spatial strain distributions
The DICE similarity coefficient (mean ± SD) between gyrencephalic and lissencephalic models for MPS distributions ranged between 0.17 ± 0.07 under the axial rotations, 0.19±0.08 under the sagittal, and 0.22 ± 0.10 under the coronal rotations, highlighting the differences between the gyrencephalic and lissencephalic models. The CSDM15 DICE ranged between 0.60 ± 0.10 under the axial, 0.52 ± 0.10 under the sagittal, and 0.59 ± 0.08 under the coronal loading (Figure 4). The DICE for CSDM15 is higher than for MPS, since CSDM15 DICE essentially discretizes the MPS distribution in steps of 1.0, while MPS DICE considers similarity in steps of 0.025.
@@
Figure 3: Peak maximum principal strain (MPS) contour plots of multiple representative subjects in the lissencephalic and the gyrencephalic models under rotational acceleration about each principal axis: a) axial, b) sagittal, and c) coronal. These MPS distributions capture the interaction of cortical folds with ventricles to result in different strain localization, especially around deeper sulci, such as the lateral (Sylvian) sulcus.
Figure 4: The similarity of the spatial distributions of strain based injury metrics in the gyrencephalic and lissencephalic model is obtained using Sørensen DICE coefficient (DICE). We find lower similarity in the MPS distributions as compared to the CSDM15 distributions under all the directions of rotations.
@@ 3.4 Effect of cortical folds on peak injury metrics
The time history plots of injury metrics, including MPS95, MPSR95, and CSDM, for all the subjects and rotation directions show that the lissencephalic models consistently under-predicted all injury metrics (Figure 5). The time to peak MPS95 and MPSR95 is also lower in the lissencephalic models for all loading directions (Figure 5), due to higher strain concentrations occurring deeper in gyrencephalic models, where the shear wave arrives later compared to the cortical surface, which are the highest strain locations in the lissencephalic models. This is also reflected in the CSDM15, which is initially higher in the lissencephalic models, before being surpassed by the gyrencephalic models Faster shear wave speed in stiffer brain tissue than the significantly more compliant CSF could also contribute to the differences in CSDM15, which rises faster in the lissencephalic model than in the gyrencephalic model.
@@
Figure 6: Bland-Altman analysis showing the mean bias and the limits of agreement (mean±1.96×SD) between the gyrencephalic (G) and lissencephalic (L) peak injury metrics: a) MPS95, b) MPSR95, and c) CSDM15. These trends show the gyrencephalic models overpredict the injury metrics compared to the lissencephalic models with high statistical significance.
@@ 3.5 Effect of modeling cortical folds on different regions of interest
3.5.1 Cerebrum
The cerebrum is geometrically different between the gyrencephalic and the lissencephalic models. The region experienced overall higher MPS95, MPSR95, and CSDM15 in the gyrencephalic than in the lissencephalic models for all three loading directions. The Bland-Altman analysis provided the bias (mean $\pm$ std. dev.) in peak MPS95 to be $0.053\pm 0.043$ (statistically significant, $p<0.005$), in peak MPSR95 to be $8.53\pm 8.56$ s-1 ($p<0.005$), and in CSDM15 to be $6.6\pm 6.4$ ($p<0.05$) (Figure 7). The bias values translate to $15.6\pm 12.7\%$, $12.7\pm 12.8\%$, and $18.2\pm 17.5\%$ of the maximum MPS95, MPSR95, and CSDM15 in the gyrencephalic models cerebrum across all models, respectively. No proportional bias was observed in any injury metric.
Since, the cerebral gray and white matter are modeled differently in the gyrencephalic and lissencephalic models (Figure 1a), the size of high-strain regions on the cortical surface can artificially cause differences in the peak injury metrics in these regions. For example, a point at a depth of 6 mm from the cortical surface will lie in the white matter in the lissencephalic model, but due to the presence of sulci 2 mm from it, it will fall in the gray matter in the gyrencephalic model. Therefore, we don’t study the gray and white matter separately, but combined.
3.5.2 Deep Brain Regions
The deep brain regions (Section 2.1) is geometrically identical between the gyrencephalic and the lissencephalic models. However, the distance between the cortical folds and the deep brain regions can be as low as $\sim$10 mm between deep sulci, such as the Sylvian fissure, or medial Cingulate sulci and outer regions of the putamen or caudate, as opposed to at least $\sim$30 mm from the cortical surface in the lissencephalic models. The region experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions ($p<0.005$), arising from the zone of strain concentrations at the base of the sulci.
We found the mean bias ($\pm$ std. dev.) in peak MPS95 was found to be $0.057\pm 0.038$, in peak MPSR95 to be $12.88\pm 7.28$ s-1, and in CSDM15 to be $17.1\pm 12.1$ (all $p<0.005$) (Figure 7). The normalized bias values are $25.3\pm 17.0\%$, $28.9\pm 16.3\%$ and $46.8\pm 33.3\%$, in MPS95, MPSR95, and CSDM15, respectively. A proportional bias was observed again in all injury metrics, with the bias increasing with the magnitude of the injury metric.
3.5.3 Corpus Callosum
The corpus callosum is geometrically the same between the gyrencephalic and the lissencephalic models, although is located close to the gyri on the medial axis. It experienced one of the highest peak values for all injury metrics and also the highest biases between the two models across all loading directions, since it sits adjacent to the lateral ventricles. The high strain in the brain tissue around the ventricles in the gyrencephalic model is passed on to the ventricles, which also experience higher strains than in the lissencephalic model. The tissues near the ventricles are stretched higher as the wave passes through the ventricles to the other side of the brain, eventually affecting the regions far removed from the cortical folds, including the corpus callosum. It experienced higher strain in the gyrencephalic than in the lissencephalic models for all three loading directions ($p<0.005$).
[…]
@@ 3.5.4 Cerebellum
The cerebellum is geometrically identical between the gyrencephalic and the lissencephalic models, and away from the cortical folds, separated by tentorium. The region experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions, due to its proximity to the fourth ventricles.
@@
The bias (mean $\pm$ std. dev.) in peak MPS95 was found to be $0.035\pm 0.042$ ($p<0.005$), in peak MPSR95 to be $7.18\pm 6.51$ s-1 ($p<0.05$), and in CSDM15 to be $10.8\pm 11.5$ ($p<0.05$) (Figure 7). The normalized bias values are $16.4\pm 19.9\%$, $16.2\pm 14.7\%$ and $29.6\pm 31.6\%$, in MPS95, MPSR95, and CSDM15, respectively. We observe proportional bias in all injury metrics, with the bias increasing with the magnitude of the injury metric.
Given the resolution of our model (1 mm), all the folds in the cerebellum are also not fully resolved, and this study keeps the cerebellum identical between the two models, and only investigates the effects of folds in the cerebrum.
3.5.5 Brain Stem
@@
The brain stem is also geometrically identical between the gyrencephalic and the lissencephalic models and separated from the cortical folds, and experienced higher peak values for all injury metrics compared to other ROIs, except the corpus callosum. The region experienced overall higher strain in the gyrencephalic models than the lissencephalic ones for all three loading directions ($\textit{p}<0.005$)due to its proximity to the fourth ventricles.
@@
The bias (mean $\pm$ std. dev.) in peak MPS95 was found to be $0.065\pm 0.050$, in peak MPSR95 to be $11.17\pm 10.85$ $s^{-1}$, and in CSDM15 to be $13.0\pm 13.7$ (all $p<0.05$) (Figure 7). The normalized bias values are $20.5\pm 15.6\%$, $17.4\pm 16.9\%$ and $35.5\pm 37.5\%$, in MPS95, MPSR95, and CSDM15, respectively. We found no proportional bias in any injury metric.
Figure 7: Different regions of interest (ROI) showing the cerebrum, cerebellum, deep brain regions, corpus callosum, and brain stem. The statistical significance of the differences is marked for the peak injury metrics MPS95, MPSR95, and CSDM15. These trends show that the gyrencephalic models predict higher peak injury metrics compared to the lissencephalic models with high statistical significance in all ROIs, except for the cerebellum, which sits far removed from the cortical folds and separated by the tentorium.
4 Discussion
In this study, we aimed to determine the need to incorporate cortical folds in FE head models for accurately predicting the risk of mTBI under head rotational acceleration events. We attempted to answer this question by comparing different mTBI injury metrics between the gyrencephalic and lissencephalic FE models of 18 subjects aged 9 - 18 years. The models were used to simulate an idealized concussive event with the rotational acceleration of 10 krad/s2 and rotational velocity of 60 rad/s about each principal axis. In this section, we discuss our results in comparison with other studies to understand the factors affecting the importance of modeling cortical folds (Section 4.1) and the limitations of our study and future works (Section 4.2).
4.1 Comparison with other studies
Several previous studies have investigated the role of cortical folds on mTBI injury metrics under head acceleration events and have found conflicting results.
Some studies are based on the full brain models, while some are based on mesoscale models of the region incorporating approximately a couple of gyri. Among the studies at the full brain scale, some predicted higher injury metrics in gyrencephalic models [14, 15], some in lissencephalic models [12], and some found different trends in different metrics [11]. The studies varied in terms of the applied loading conditions, the analyzed injury metrics, and the material properties used in the models. While the studies focused on impact loading, most studies only incorporated translational accelerations [15, 13, 11, 14]. It has been shown that during an impact, the resulting translational acceleration leads primarily to high pressure, while the angular kinematics primarily leads to the shear strains in the brain tissue [47]. The low pressure from translational acceleration is correlated to cavitation in the brain tissue, while shear strain is directly related to neuronal stretch and death. Since an impact loading results in negative pressures much lower than cavitation thresholds, the rotational acceleration of the head is more critical in an impact loading [48, 30]. Therefore, our study focuses on rotational head accelerations, and we can only compare our results with studies that incorporate angular acceleration.
The low impact of translational acceleration on brain tissue is reflected in the low strains reported in these studies ($5\%$). Fagan et al. [15] found higher strains in gyrencephalic models but similar pressure response in the two models; Mazurkiewicz et al. [14] also found higher strains in gyrencephalic model; Saez et al. [13] found similar strains in the two models but lower pressures and maximum principal stresses in gyrencephalic model; and Song et al. [11] found no difference in the pressure, higher stress and lower strains in gyrencephalic models. The differences in these results could stem from the differences in the models used. Saez et al. modeled a 3D models of a human (high gyrification), a macaque (lower gyrification), and a mouse (lissencephalic) scaled to the same size, Fagan et al. developed a 2D gyrencephalic and lissencephalic models of human axial slice, Mazurkiewicz et al. conducted experiments on a 2D gyrencephalic and lissencephalic surrogates based on coronal slice of porcine brain, and Song et al. studied the sagittal slice of a human head. The studies also differed in terms of material properties, e.g. Fagan et al. modeled the CSF as an Eulerian mesh, resulting in separation between the brain and the CSF; while the CSF material modeling was not reported in the Saez et al. paper. These differences make the direct comparison difficult to identify the factors affecting the role of cortical folds.